OpenStudy (anonymous):

Is the line through points P(-8, -10) and Q(-5, -12) perpendicular to the line through points R(9, -6) and S(17, -5)? Promised medal to helpful answer! c:

OpenStudy (anonymous):

Can you find the slope of both pairs of lines?

OpenStudy (anonymous):

@Ookami47

OpenStudy (anonymous):

Hm. so ((-10) - )-12)) = -21 would equal the run right? ._. and the rise would be ((-8_ - (-5) = -13?

OpenStudy (anonymous):

Would -13/-21 be one of the slopes? I'm just absolutely horrible when it comes to slopes..

OpenStudy (anonymous):

no these lines arenot perpendicular because product of slope of these two lines should be -1means slope of PQ*slope of RS=-1

OpenStudy (anonymous):

Yes that is correct. So the slope would be 13/21 right?

OpenStudy (anonymous):

Now find the other slope.

OpenStudy (anonymous):

@CUTE_AKKI : I was getting to that but I wanted him to do out the work :P .

OpenStudy (anonymous):

-8/11? o:

OpenStudy (anonymous):

Or am I backwards.. xD

OpenStudy (anonymous):

Should be 8/1 or just 8 :) .

OpenStudy (anonymous):

Oh. cx alrighty. What comes next?

OpenStudy (anonymous):

Two lines are said to be Perpendicular if the slopes of both lines are negative reciprocals of each other.

OpenStudy (anonymous):

As @CUTE_AKKI said.

OpenStudy (anonymous):

Clearly, you can see they are not.

OpenStudy (anonymous):

Obviously c:

OpenStudy (anonymous):

So therefore, they are NOT Perpendicular :P .

OpenStudy (anonymous):

Sweet! So I would just answer that they wouldn't be perpendicular due to the slopes not being negative reciprocals of each other?

OpenStudy (anonymous):

ok....let m1 be slope of PQ =(-12-(-10))/(-5-(-8))=-2/3=m1 let m2 be slope of RS=((-5-(-6))/(17-9)=1/8=m2 now m1*m2 is not = -1 so given lines aren't perpendicular

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

I see... thank you for your help you guys. c: